Show how the average energy in a Fermi gas can be expressed in terms of Ef0 


1
Expert's answer
2021-01-20T07:24:21-0500

When we have the 1-D box, then the energy of the electron in the box can be represents as below


"E_n=E_0+\frac{ℏ^2π^2}{2mL^2}n^2"


We know that Fermi level is the energy of the last electron, if we have 2N electron then we can find the mean energy of the electron with the given equation,


"E_{av}=E_0+\frac{2\Sigma_1^NE_n}{2N}"


"\Rightarrow E_{av}=E_0+\frac{2\Sigma_1^N E_o +\frac{ℏ^2π^2}{2mL^2}n^2}{2N}"


we know that "\Sigma n^2=\frac{N(N+1)(2N+1)}{6}=\frac{N^3}{3}+\frac{N^2}{2}+\frac{N}{6}"


If we consider N is large value, then we can ignore the second and third terms.


If we assume that there is lots of electron, then "E_{av}=E_o+\frac{ℏ^2π^2 N^3}{2mL^23N}"


But, Fermi level "E_f=E_o+\frac{ℏ^2π^2 }{2mL^2}N^2"


But if "E_o=0"


"E_{av}=\frac{E_f}{3}"